OpenStudy (anonymous):

WILL MEDAL AND FAN BEST ANSWER!! I NEED THE ANSWER -42 + (-49) = ? AND -36 + 55 =?

OpenStudy (haseeb96):

-42 -49 = -91 and 55 -36 =19

OpenStudy (anonymous):

-42 -49 = -91 and -36+55 =19

OpenStudy (anonymous):

thanks